You can invest in a bank account that has an APR of 8%. You want to make equal deposits every semester (six months) of $500 for the next 5 years. How much will you have accumulated at the end of the 5 years?

A. $14, 889.04

B. 6,003.05

C. $7,243.28

D, $6,795.16
Answer #1

Future value of Annuity = A [((1+r)n-1) / r]

Where

A - Annuity payment = 500

r - rate per period = .08/2 = .04

n - no. of periods = 5*2 = 10

Future value of Annuity = 500* [(1.04^10-1) / .04]

= 500* [(1.48024428492-1) / .04]

= 500*12.006107123

= 6003.05